OpenStudy (anonymous):

Determine the x- and y- intercepts for the graph defined by the given equation. y = x + 8 a. x-intercept is (0, 8) y-intercept is ( -8, 0) c. x-intercept is ( -8, 0) y-intercept is (0, 8) b. x-intercept is (0, -8) y-intercept is ( 8, 0) d. x-intercept is ( 8, 0) y-intercept is (0, -8)

Nnesha (nnesha):

y =mx + b here b is y-intercept y-intercept cross y-axis when x=0 so what is b in your equation ??

OpenStudy (anonymous):

8?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

and x -intercept when line cross x-axis and y = 0 so use ur equation y = x+8 replace y by 0 and solve for x

OpenStudy (anonymous):

0=x+8 -8=x

Nnesha (nnesha):

yesh that's right

OpenStudy (anonymous):

(-8, 0) ?? c?

Nnesha (nnesha):

yes c

OpenStudy (anonymous):

Thank you

Nnesha (nnesha):

my pleasure :)